OpenStudy (maddi0453):

Which function in vertex form is equivalent to f(x) = 4 + x2 – 2x? f(x) = (x – 1)2 + 3 f(x) = (x – 1)2 + 5 f(x) = (x + 1)2 + 3 f(x) = (x + 1)2 + 5

OpenStudy (maddi0453):

@nightowl

OpenStudy (mathmale):

Please expres exponentiation properly: Correct: f(x) = 4 + x^2 – 2x Vague: f(x) = 4 + x2 – 2x

OpenStudy (mathmale):

To do this problem, you need to re-write the given equation in the vertex form by "completing the square." Notice that f(x)=4+x^2-2x can be rewritten as f(x)=x^2 - 2x - 4

OpenStudy (mathmale):

Re-write x^2 - 3x as a perfect square.. If neessary, read up on "completing the square."